# Uncovering Vulnerabilities of LLM-Assisted Cyber Threat Intelligence

### v2

@@ B.1 Used Benchmark in Evaluation
Table LABEL:tab:src-bench-stat presents the original scales of the datasets used in our evaluation.
@@
CTIBench (Alam et al., 2024) Cyber Threat Intelligence 5,610 5 N/A
SevenLLM-Bench (Ji et al., 2024) Report Analyzing 92,701 28 N/A
SWE-Bench (Jimenez et al., 2023) Bug fixing 2,294 12 1
CYBERTEAM (Liu et al., 2025) Blue-team threat hunting 452,293 30 23
@@ B.2 Used Real-World Databases and Platforms in Evaluation
National Vulnerability Database (NVD). The NVD (National Institute of Standards and Technology (2024), NIST) serves as the canonical repository for software vulnerabilities and their CVSS severity scores. We utilize NVD entries to support all ❶–❹ tasks. Specifically, we map vulnerabilities in benchmark items to NVD records in order to standardize CVE identifiers, extract official CVSS vector strings, and obtain temporal metadata (publication and modification dates). These fields allow us to align system-environment observations with ground-truth vulnerability characteristics, and to test forecasting models that predict exploit likelihood (e.g., by contrasting NVD base scores against EPSS-derived estimates).
@@
Exploit Prediction Scoring System (EPSS). EPSS (Jacobs & Romanosky, 2021) provides probabilistic estimates of the likelihood that a given CVE will be exploited in the wild. We use EPSS scores directly for ❸ prediction tasks, both as a source of labels (ground-truth exploitation likelihood) and as a reference distribution to evaluate calibration of LLM-based forecasts. EPSS time-series updates also enable temporal correlation experiments, where we assess whether models capture shifts in exploitation likelihood following disclosure, patch release, or threat-actor reuse.
@@
MITRE ATT&CK and CAPEC. The ATT&CK (mit, 2020) knowledge base encodes adversarial tactics, techniques, and procedures (TTPs) in a structured taxonomy. We map CTI tasks involving entity extraction, campaign attribution, and mitigation alignment to ATT&CK entries. For instance, when evaluating ❷ attribution, extracted TTPs from LLM outputs are compared against ATT&CK technique identifiers to assess correctness. We also leverage CAPEC (Common Attack Pattern Enumeration and Classification) (MITRE Corporation, 2024a) to validate abstract attack patterns referenced in benchmark items, particularly for mapping contextualized logs or IOCs to higher-level adversarial behaviors.
@@
MISP (Malware Information Sharing Platform). MISP (MISP Project, 2011) serves as a community-driven threat intelligence sharing platform, containing structured feeds of indicators of compromise (IOCs), malware samples, and infrastructure metadata. We use MISP to enrich contextualization tasks (❶) by grounding benchmark instances in realistic IOC–malware–actor relationships. For example, when a dataset item involves resolving a suspicious domain, we verify its presence in MISP feeds and align it to associated threat actors or malware families. This enrichment supports evaluation of LLMs’ ability to normalize IOCs and link them to campaigns.
@@
VirusTotal. VirusTotal (Google LLC, 2004) aggregates antivirus detections and malware analysis results across a large corpus of submitted files, domains, and URLs. We leverage VirusTotal reports for ❶ contextualization and ❷ attribution tasks. In contextualization, VirusTotal tags (e.g., malware family labels, sandbox behavior summaries) serve as auxiliary ground-truth for tasks like malware family mapping. In attribution, we analyze infrastructure overlap by checking whether related domains or IPs have been co-reported in VirusTotal samples, allowing us to validate LLM predictions about infrastructure reuse.
@@
Open Threat Exchange (OTX). AlienVault’s OTX (AlienVault, 2012) provides community-curated threat pulses (collections of IOCs associated with specific campaigns or malware). We use OTX primarily for ❷ attribution and ❸ prediction: pulses give us labeled groupings of IOCs tied to campaigns, which we then cross-check against LLM-predicted campaign attributions.
@@
Security Advisories and Vendor Bulletins. Finally, vendor advisories (e.g., Microsoft, Cisco, Progress Software) and public CERT bulletins provide authoritative patching and mitigation recommendations (Multiple Vendors (2000), Microsoft, Cisco, Progress Software, etc.). We incorporate these resources into ❹ mitigation tasks by aligning recommended countermeasures with benchmark items. For instance, in patch recommendation evaluation, the correct answer set is derived from vendor bulletins rather than from secondary threat reports. Similarly, YARA and Sigma rule examples are drawn from advisory-linked repositories, ensuring that response summarization tasks are grounded in practical remediation steps.
@@ B.4 Data Statistics
Affected Systems 38.6K
Attack Infrastructure 42.7K
Vulnerability Linking 35.2K
Malware Family Mapping 29.8K
IOC Normalization 33.1K
Threat Report Alignment 26.4K
Event Timeline Construction 19.7K
Graph Population 22.9K
Source Reliability Scoring 16.2K
@@
Threat Actor Linking 21.6K
TTP Extraction 30.4K
Campaign Attribution 17.9K
Infrastructure Reuse 19.8K
Language/Style Profiling 12.7K
False Flag Detection 9.4K
Evidence Weighting 15.2K
Relation Graph Building 18.4K
@@
Exploit Likelihood 24.6K
Impact Forecast 14.8K
Target Sector Prediction 13.2K
Campaign Escalation 11.7K
@@
Patch Recommendation 19.3K
Rule Generation (YARA) 12.4K
Response Summarization 16.8K
Mitigation–TTP Mapping 14.7K
Defensive Playbook Gen 11.6K
Countermeasure Ranking 13.1K
Incident Ticket Generation 12.2K
@@ B.5 Model Name and Version
Here we detail the used LLMs in our extensive evaluation (§2.2), corresponding to the abbrevated names used in Table 2 and in Section §4: G5–GPT-5, Go4–GPT-o4 mini, CLD–Claude Sonnet 4, GEM–Gemini 2.5, LL70–Llama-3.1-70B-Instruct, MIX–Mixtral-8x7B-Instruct-v0.1, QWN–Qwen2.5-14B-Instruct, GRK–Grok-2, FSC–Foundation-Sec-8B (Weerawardhena et al., 2025), CB0–Cyber-Zero (Zhuo et al., 2025), ZYS–ZySec-AI-SecurityLLM (ZySec AI, 2024), LLY–Lily-Cybersecurity-7B-v0.2 (Labs, 2024), CBS–CyberBase-13b (CyberNative, 2024), SPT–clouditera-secgpt (clouditera, 2025), DHT–DeepHat-V1-7B (DeepHat, 2025).
@@ Appendix C Complementary Details for §3
C.1 Algorithm
This part presents algorithm for our failure categorization methods:
Input:
Original dataset $\mathcal{D}_{o}=\{x_{i}\}_{i=1}^{N}$ (CTI instances);
Predicted reports $\{y_{i}\}$ (model-generated outputs);
Ground-truth reports $\{r_{i}\}$ (reference advisories);
Evaluation metric $Sim(\cdot,\cdot)$ (e.g., BLEU);
Quantile step $\delta=0.05$;
1
Output:
Stratified failure group $\mathcal{D}$;
2
3foreach $x_{i}\in\mathcal{D}_{o}$ do
4   Compute similarity score $s_{i}=Sim(y_{i},r_{i})$;
5
6 end foreach
7
8Sort $\{s_{i}\}$ in ascending order;
9 Partition into candidate strata $\mathcal{G}_{\delta}$ of size $\delta$ (quantile bins);
10
// Step 1: Establish anchors
11 Manually inspect representative samples in selected $\mathcal{G}_{\delta}$;
12 Record score ranges for “correct” vs. “failed” anchors;
13
// Step 2: Assign remaining instances
14 foreach $x_{i}$ not manually inspected do
15 if $s_{i}$ within failed anchor range then
    $x_{i}\rightarrow\mathcal{D}$ ;  // Label $x_{i}$ as failure
16
17 else
18    if $s_{i}$ within correct anchor range then
19         Label $x_{i}$ as non-failure (i.e., correct);
20
21    else
22         Manually inspect $x_{i}$ (boundary case);
23
24      end if
25
26   end if
27
28 end foreach
29
// Step 3: Termination condition
30 Stop once no new failure modes appear and distribution of failure modes stabilizes across strata;
31
return Failure group $\mathcal{D}$;
Algorithm 1 Stratification for Failure Analysis (RQ1)
@@ 30 Stop once no new failure modes appear and distribution of failure modes stabilizes across strata;
2 Human annotators $H$ inspect a subset $\mathcal{D}_{0}\subset\mathcal{D}$;
3 Derive initial taxonomy $\mathcal{T}_{0}=\{t_{1},\dots,t_{k}\}$;
4
@@ 3 Derive initial taxonomy $\mathcal{T}_{0}=\{t_{1},\dots,t_{k}\}$;
5 foreach $x_{j}\in\mathcal{D}\setminus\mathcal{D}_{0}$ do
6   Assign label $y_{j}\in\mathcal{T}_{m}\cup\{\texttt{other}\}$ using $f_{\theta}$;
7
8 end foreach
9
@@ 6   Assign label $y_{j}\in\mathcal{T}_{m}\cup\{\texttt{other}\}$ using $f_{\theta}$;
10 Collect $\mathcal{O}_{m}=\{x_{j}\mid y_{j}=\texttt{other}\}$;
11 Human annotators $H$ inspect $\mathcal{O}_{m}$;
12 If new modes $\Delta\mathcal{T}$ found, update taxonomy:
13 $\mathcal{T}_{m+1}\leftarrow\mathcal{T}_{m}\cup\Delta\mathcal{T}$;
14
@@ 13 $\mathcal{T}_{m+1}\leftarrow\mathcal{T}_{m}\cup\Delta\mathcal{T}$;
15 Repeat Steps 2–3 until $\Delta\mathcal{T}=\varnothing$ and coverage $\geq\rho$;
16 Set $\mathcal{T}^{*}=\lim_{m\to\infty}\mathcal{T}_{m}$;
17
18return stabilized taxonomy $\mathcal{T}^{*}$;
@@ 16 Set $\mathcal{T}^{*}=\lim_{m\to\infty}\mathcal{T}_{m}$;
2 foreach $x_{i}\in\mathcal{D}$ do
3   Each model $a\in\Theta$ independently assigns $\hat{y}_{i}^{(1,a)}\in\mathcal{T}$;
4
5 end foreach
6
@@ 3   Each model $a\in\Theta$ independently assigns $\hat{y}_{i}^{(1,a)}\in\mathcal{T}$;
7 foreach $x_{i}\in\mathcal{D}$ do
8   Each model $a\in\Theta$ observes $\{\hat{y}_{i}^{(1,b)}\}_{b\in\Theta}$;
9   Refine reasoning and output $\hat{y}_{i}^{(2,a)}$;
10
11 end foreach
12
@@ 9   Refine reasoning and output $\hat{y}_{i}^{(2,a)}$;
13 Define uncertain set:
14 $\mathcal{U}=\{x_{i}\mid\exists a:\text{Var}(\{\hat{y}_{i}^{(a)}\}_{a\in\Theta})>0\;\;\textbf{or}\;\;\hat{y}_{i}^{(1,a)}\neq\hat{y}_{i}^{(2,a)}\}$;
15
16Human annotators $H$ inspect all $x_{i}\in\mathcal{U}$ and assign final labels;
17
@@ 13 Define uncertain set:
18 Instances $\mathcal{D}\setminus\mathcal{U}$ take majority-agreed labels from models;
19 Instances $\mathcal{U}$ take human-verified labels;
20
21return categorized dataset with reliable failure mode assignments;
@@ 19 Instances $\mathcal{U}$ take human-verified labels;
C.2 Determining Vulnerability Types
@@ D.1 Root Causes of Vulnerabilities (Continue to RQ5)
Case Study. As reported by Group-IB (Group-IB, 2025), analysts observed infrastructure overlap in domains or servers co-mentioned across multiple campaigns attributed to MuddyWater. However, further investigation revealed that although the same infrastructure was used or appeared in reports, the operational characteristics (targeting, malware payloads, attack vectors) differed significantly between those campaigns. Because relation extraction or graph building techniques often rely on co-occurrence of infrastructure to infer actor reuse, models might erroneously link distinct campaigns to the same threat actor, or assume a shared campaign lineage, solely based on the shared infrastructure.
@@
Case Study. In Microsoft’s June 2025 Patch Tuesday advisory, 65 CVEs were addressed, including a zero-day vulnerability exploited in the wild. The advisory lists patches for many components including Microsoft Office, Visual Studio, Windows Kernel, WebDAV, SMB, and others (Research, 2025). Because patching is broad and multiple fixes are mentioned in the same bulletin, a CTI model might treat all listed patches as equally urgent defense actions — even though only a small subset correspond to vulnerabilities currently exploited. This leads organizations or systems to over-allocate resources toward non-critical patching, or generate mitigation playbooks that include countermeasures not immediately relevant to active threats.
@@
Case Study. Multiple CVEs share many IoCs (e.g., IP addresses and domain names) across different CTI provider feeds (Kodituwakku et al., 2023). Over time, some of these IoCs appear repeatedly in contexts of various vulnerabilities, even though not all of them are actually exploited in relation to each CVE. Because prediction models often take frequency of IoC appearance as a strong signal, they tend to assign higher risk to these CVEs or predict escalation based on the reused IoCs. In this way, reuse of IPs/domains (which may simply reflect broad surveillance coverage or shared infrastructure, not genuine exploit activity) inflates perceived future threat likelihoods.
@@
Case Study. A concrete example is the evolution of exploit activity for CVE‑2020‑1472 (Zerologon). Initially, after its disclosure, exploit usage was limited and many datasets treated the vulnerability as low risk. Over time, threat actors began leveraging Zerologon widely, including in ransomware and lateral movement toolchains. EclecticIQ’s long‑term analysis demonstrates how risk associated with this vulnerability evolved substantially across months, with early scarce activity later giving way to broad exploit adoption (Team, 2022). Predictive models that do not de‑emphasize earlier false negatives risk underestimating exploitability or generating contradictory forecasts. This case underscores how temporal contradiction — mixing outdated low‑risk labeling with current high exploitation data — can destabilize predictions of vulnerability exploit likelihood.
@@
Conflicting report (2.2) in ❶ contextualization. Conflicting reports frequently arise during contextualization because raw observations are aggregated from diverse sources — vendor advisories, community threat feeds, and open reports — that describe incidents with varying detail and emphasis. For example, one feed may assert that a suspicious log entry corresponds to exploitation of a particular CVE, while another links the same log to an entirely different vulnerability or malware family. Topic modeling and event extraction pipelines, designed to cluster narratives and structure incidents, may then merge these contradictory claims into a single enriched context. This creates confusion over which CVE, TTP, or malware family is actually relevant, with the root cause being the absence of a standardized ground truth and the reliance on partially overlapping but inconsistent external reports.
@@
Semantic conflict (2.3) in ❷ attribution (Why not influential). In attribution, semantic conflict is less influential because attribution reasoning does not hinge primarily on the naming of entities but on relational patterns and behavioral signatures. For instance, identifying an actor depends more on shared infrastructure reuse, campaign TTPs, or linguistic style than on whether a malware sample is labeled “PlugX” or “Korplug.” Relation extraction and event graph construction are designed to capture structural patterns rather than surface semantics, allowing attribution models to tolerate naming inconsistencies so long as the underlying features remain consistent. While semantic divergence may create minor noise, it rarely changes the actor assignment itself; two labels referring to the same malware family still connect to the same infrastructure nodes or TTPs. Thus, attribution tasks are comparatively robust to semantic conflicts.
@@
Case Study. In our attribution dataset, we observed a campaign exploiting CVE‑2022‑30190 (Follina Microsoft Office Remote Code Execution) whose behavioral signature combined TTPs typical of both Group X and Group Y (e.g. custom VBA macro infection + unusual DNS tunneling), yet also introduced a new lateral movement module not seen before. The attribution model forced the campaign into Group Y because the overlapping macro artifacts and DNS patterns had been heavily associated with Group Y in training. However, deeper analysis of payload internals and command semantics showed the lateral movement logic was markedly different from any known Group Y campaign and the targeting region also diverged. Because the attribution system attempted to ”fit” the new pattern into the nearest known actor schema rather than flagging it as novel, the campaign was misattributed. This case illustrates how unseen pattern adoption can lead attribution models to overcommit to the nearest known archetype, thereby misclassifying new or hybrid campaigns.

### v3

@@ B.1 Used Benchmark in Evaluation
Table 8 presents the original scales of the datasets used in our evaluation.
@@
CTIBench [alam2024ctibench] Cyber Threat Intelligence 5,610 5 N/A
SevenLLM-Bench [ji2024sevenllm] Report Analyzing 92,701 28 N/A
SWE-Bench [jimenez2023swe] Bug fixing 2,294 12 1
CYBERTEAM [liu2025benchmarking] Blue-team threat hunting 452,293 30 23
@@ B.2 Used Real-World Databases and Platforms in Evaluation
National Vulnerability Database (NVD). The NVD [nvd] serves as the canonical repository for software vulnerabilities and their CVSS severity scores. We utilize NVD entries to support all ❶–❹ tasks. Specifically, we map vulnerabilities in benchmark items to NVD records in order to standardize CVE identifiers, extract official CVSS vector strings, and obtain temporal metadata (publication and modification dates). These fields allow us to align system-environment observations with ground-truth vulnerability characteristics, and to test forecasting models that predict exploit likelihood (e.g., by contrasting NVD base scores against EPSS-derived estimates).
@@
Exploit Prediction Scoring System (EPSS). EPSS [Jacobs2021EPSS] provides probabilistic estimates of the likelihood that a given CVE will be exploited in the wild. We use EPSS scores directly for ❸ prediction tasks, both as a source of labels (ground-truth exploitation likelihood) and as a reference distribution to evaluate calibration of LLM-based forecasts. EPSS time-series updates also enable temporal correlation experiments, where we assess whether models capture shifts in exploitation likelihood following disclosure, patch release, or threat-actor reuse.
@@
MITRE ATT&CK and CAPEC. The ATT&CK [mitreattack] knowledge base encodes adversarial tactics, techniques, and procedures (TTPs) in a structured taxonomy. We map CTI tasks involving entity extraction, campaign attribution, and mitigation alignment to ATT&CK entries. For instance, when evaluating ❷ attribution, extracted TTPs from LLM outputs are compared against ATT&CK technique identifiers to assess correctness. We also leverage CAPEC (Common Attack Pattern Enumeration and Classification) [capec] to validate abstract attack patterns referenced in benchmark items, particularly for mapping contextualized logs or IOCs to higher-level adversarial behaviors.
@@
MISP (Malware Information Sharing Platform). MISP [MISP] serves as a community-driven threat intelligence sharing platform, containing structured feeds of indicators of compromise (IOCs), malware samples, and infrastructure metadata. We use MISP to enrich contextualization tasks (❶) by grounding benchmark instances in realistic IOC–malware–actor relationships. For example, when a dataset item involves resolving a suspicious domain, we verify its presence in MISP feeds and align it to associated threat actors or malware families. This enrichment supports evaluation of LLMs’ ability to normalize IOCs and link them to campaigns.
@@
VirusTotal. VirusTotal [VirusTotal] aggregates antivirus detections and malware analysis results across a large corpus of submitted files, domains, and URLs. We leverage VirusTotal reports for ❶ contextualization and ❷ attribution tasks. In contextualization, VirusTotal tags (e.g., malware family labels, sandbox behavior summaries) serve as auxiliary ground-truth for tasks like malware family mapping. In attribution, we analyze infrastructure overlap by checking whether related domains or IPs have been co-reported in VirusTotal samples, allowing us to validate LLM predictions about infrastructure reuse.
@@
Open Threat Exchange (OTX). AlienVault’s OTX [OTX] provides community-curated threat pulses (collections of IOCs associated with specific campaigns or malware). We use OTX primarily for ❷ attribution and ❸ prediction: pulses give us labeled groupings of IOCs tied to campaigns, which we then cross-check against LLM-predicted campaign attributions.
@@
Security Advisories and Vendor Bulletins. Finally, vendor advisories (e.g., Microsoft, Cisco, Progress Software) and public CERT bulletins provide authoritative patching and mitigation recommendations [VendorAdvisories]. We incorporate these resources into ❹ mitigation tasks by aligning recommended countermeasures with benchmark items. For instance, in patch recommendation evaluation, the correct answer set is derived from vendor bulletins rather than from secondary threat reports. Similarly, YARA and Sigma rule examples are drawn from advisory-linked repositories, ensuring that response summarization tasks are grounded in practical remediation steps.
@@ B.4 Data Statistics
Affected Systems 3.9K
Attack Infrastructure 4.1K
Vulnerability Linking 3.5K
Malware Family Mapping 2.1K
IOC Normalization 3.1K
Threat Report Alignment 2.0K
Event Timeline Construction 1.5K
Graph Population 2.3K
Source Reliability Scoring 1.6K
@@
Threat Actor Linking 2.1K
TTP Extraction 3.0K
Campaign Attribution 1.8K
Infrastructure Reuse 1.2K
Language/Style Profiling 1.3K
False Flag Detection 0.9K
Evidence Weighting 1.5K
Relation Graph Building 1.8K
@@
Exploit Likelihood 2.5K
Impact Forecast 1.5K
Target Sector Prediction 1.3K
Campaign Escalation 1.1K
@@
Patch Recommendation 1.9K
Rule Generation (YARA) 1.2K
Response Summarization 1.8K
Mitigation–TTP Mapping 1.5K
Defensive Playbook Gen 1.1K
Countermeasure Ranking 1.3K
Incident Ticket Generation 1.2K
@@ B.5 Model Name and Version
Here we detail the used LLMs in our extensive evaluation (§2.2), corresponding to the abbreviated names used in Table 2 and in Section §4: G5–GPT-5, Go4–GPT-o4 mini, CLD–Claude Sonnet 4, GEM–Gemini 2.5, LL70–Llama-3.1-70B-Instruct, MIX–Mixtral-8x7B-Instruct-v0.1, QWN–Qwen2.5-14B-Instruct, GRK–Grok-2, FSC–Foundation-Sec-8B [weerawardhena2025llama], CB0–Cyber-Zero [zhuo2025cyber], ZYS–ZySec-AI-SecurityLLM [ZySecAI], LLY–Lily-Cybersecurity-7B-v0.2 [SegoLily], CBS–CyberBase-13b [CyberNative], SPT–clouditera-secgpt [SecGPT], DHT–DeepHat-V1-7B [DeepHat].
@@ Appendix C Complementary Details for §3
Human annotators $H$ inspect a subset $\mathcal{D}_{0}\subset\mathcal{D}$
Derive initial taxonomy $\mathcal{T}_{0}=\{t_{1},\dots,t_{k}\}$
@@
foreach $x_{j}\in\mathcal{D}\setminus\mathcal{D}_{0}$ do
  Assign label $y_{j}\in\mathcal{T}_{m}\cup\{\texttt{other}\}$ using $f_{\theta}$
end foreach
@@
Collect $\mathcal{O}_{m}=\{x_{j}\mid y_{j}=\texttt{other}\}$
Human annotators $H$ inspect $\mathcal{O}_{m}$
If new modes $\Delta\mathcal{T}$ found, update taxonomy:
$\mathcal{T}_{m+1}\leftarrow\mathcal{T}_{m}\cup\Delta\mathcal{T}$
@@
Repeat Steps 2–3 until $\Delta\mathcal{T}=\varnothing$ and coverage $\geq\rho$
Set $\mathcal{T}^{*}=\lim_{m\to\infty}\mathcal{T}_{m}$
return stabilized taxonomy $\mathcal{T}^{*}$
@@
foreach $x_{i}\in\mathcal{D}$ do
  Each model $a\in\Theta$ independently assigns $\hat{y}_{i}^{(1,a)}\in\mathcal{T}$
end foreach
@@
foreach $x_{i}\in\mathcal{D}$ do
  Each model $a\in\Theta$ observes $\{\hat{y}_{i}^{(1,b)}\}_{b\in\Theta}$
  Refine reasoning and output $\hat{y}_{i}^{(2,a)}$
end foreach
@@
Define uncertain set:
$\mathcal{U}=\{x_{i}\mid\exists a:\text{Var}(\{\hat{y}_{i}^{(a)}\}_{a\in\Theta})>0\;\;\textbf{or}\;\;\hat{y}_{i}^{(1,a)}\neq\hat{y}_{i}^{(2,a)}\}$
Human annotators $H$ inspect all $x_{i}\in\mathcal{U}$ and assign final labels
@@
Instances $\mathcal{D}\setminus\mathcal{U}$ take majority-agreed labels from models
Instances $\mathcal{U}$ take human-verified labels
return categorized dataset with reliable failure mode assignments
@@
C.1 Determining Vulnerability Types
@@ D.1 Root Causes of Vulnerabilities (Continue to RQ5)
Case Study. As reported by Group-IB [groupib2025muddywater], analysts observed infrastructure overlap in domains or servers co-mentioned across multiple campaigns attributed to MuddyWater. However, further investigation revealed that although the same infrastructure was used or appeared in reports, the operational characteristics (targeting, malware payloads, attack vectors) differed significantly between those campaigns. Because relation extraction or graph building techniques often rely on co-occurrence of infrastructure to infer actor reuse, models might erroneously link distinct campaigns to the same threat actor, or assume a shared campaign lineage, solely based on the shared infrastructure.
@@
Case Study. In Microsoft’s June 2025 Patch Tuesday advisory, 65 CVEs were addressed, including a zero-day vulnerability exploited in the wild. The advisory lists patches for many components including Microsoft Office, Visual Studio, Windows Kernel, WebDAV, SMB, and others [tenable2025junepatchtuesday]. Because patching is broad and multiple fixes are mentioned in the same bulletin, a CTI model might treat all listed patches as equally urgent defense actions — even though only a small subset correspond to vulnerabilities currently exploited. This leads organizations or systems to over-allocate resources toward non-critical patching, or generate mitigation playbooks that include countermeasures not immediately relevant to active threats.
@@
Case Study. Multiple CVEs share many IoCs (e.g., IP addresses and domain names) across different CTI provider feeds [kodituwakku2023temporal]. Over time, some of these IoCs appear repeatedly in contexts of various vulnerabilities, even though not all of them are actually exploited in relation to each CVE. Because prediction models often take frequency of IoC appearance as a strong signal, they tend to assign higher risk to these CVEs or predict escalation based on the reused IoCs. In this way, reuse of IPs/domains (which may simply reflect broad surveillance coverage or shared infrastructure, not genuine exploit activity) inflates perceived future threat likelihoods.
@@
Case Study. A concrete example is the evolution of exploit activity for CVE‑2020‑1472 (Zerologon). Initially, after its disclosure, exploit usage was limited and many datasets treated the vulnerability as low risk. Over time, threat actors began leveraging Zerologon widely, including in ransomware and lateral movement toolchains. EclecticIQ’s long‑term analysis demonstrates how risk associated with this vulnerability evolved substantially across months, with early scarce activity later giving way to broad exploit adoption [eclecticiq2022zerologon]. Predictive models that do not de‑emphasize earlier false negatives risk underestimating exploitability or generating contradictory forecasts. This case underscores how temporal contradiction — mixing outdated low‑risk labeling with current high exploitation data — can destabilize predictions of vulnerability exploit likelihood.
@@
Conflicting report (2.2) in ❶ contextualization. Conflicting reports frequently arise during contextualization because raw observations are aggregated from diverse sources (e.g., vendor advisories, community threat feeds, and open reports) that describe incidents with varying detail and emphasis. For example, one feed may assert that a suspicious log entry corresponds to exploitation of a particular CVE, while another links the same log to an entirely different vulnerability or malware family. Topic modeling and event extraction pipelines, designed to cluster narratives and structure incidents, may then merge these contradictory claims into a single enriched context. This creates confusion over which CVE, TTP, or malware family is actually relevant, with the root cause being the absence of a standardized ground truth and the reliance on partially overlapping but inconsistent external reports.
@@
Case Study. In our attribution dataset, we observed a campaign exploiting CVE‑2022‑30190 (Follina Microsoft Office Remote Code Execution) whose behavioral signature combined TTPs typical of both Group X and Group Y (e.g. custom VBA macro infection + unusual DNS tunneling), yet also introduced a new lateral movement module not seen before. The attribution model forced the campaign into Group Y because the overlapping macro artifacts and DNS patterns had been heavily associated with Group Y in training. However, deeper analysis of payload internals and command semantics showed the lateral movement logic was markedly different from any known Group Y campaign and the targeting region also diverged. Because the attribution system attempted to "fit" the new pattern into the nearest known actor schema rather than flagging it as novel, the campaign was misattributed. This case illustrates how unseen pattern adoption can lead attribution models to overcommit to the nearest known archetype, thereby misclassifying new or hybrid campaigns.